OpenStudy (anonymous):

Solve the equation -7(t-6)-(t+7)=10

OpenStudy (anonymous):

t=39/8

OpenStudy (anonymous):

WAIT, ITS WRONG!!!!

OpenStudy (anonymous):

t=35/8

OpenStudy (anonymous):

WAIT ITS STILL WRONG!!!

OpenStudy (anonymous):

OK, I got it this time... t=25/8

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

hehehe... i think i'm losing concentration or something... the mothballs in my brain are coming back to haunt me...